OpenStudy (anonymous):

-10=-2+1/2d solve for d.

OpenStudy (anonymous):

-8=1/2d d=-16

OpenStudy (anonymous):

-16

OpenStudy (lgbasallote):

is it... \[\large -10 = -2 + \frac {1}{2d} \] or \[\large -10 = -2 + \frac {1}{2}d \]

OpenStudy (anonymous):

the 2nd 1